on taking place in a Given the information below, calculate the ΔH for a reacti coffee cup calorimeter in which 175 mL water changes temperature 31 °C 3) from 13 Cto density (H20)-0.998 g/mL specific heat (s) = 4.1841/g, K
Answer #1

Answer

\DeltaH = -30.69kJ

Explanation

mass of water = 175ml * 0.998g/ml = 174.65g

temperature difference , \DeltaT = 73 - 31 = 42 degree C

Heat absorbed by water , q = 174.65g *42degree C * 4.184J/g degree C =30.69kJ

\DeltaH = -q

= - 30.69kJ